OpenStudy (anonymous):

Make a 5% offer on a Volkswagen New Beetle that has a sticker price of $22,120 and a dealer cost of $20,856. (Round your answer to the nearest dollar.)

OpenStudy (anonymous):

A club with 29 members is to select five officers (president, vice president, secretary, treasurer, and historian). In how many ways can this be done?

OpenStudy (shane_b):

In the future please post the questions separately to avoid confusion when people start replying. For the first question, I assume a "5% offer" is 5% above dealer cost. If that's the case it would be:\[$20,856*1.05=?\]

OpenStudy (shane_b):

For the second one I believe it would be:\[\frac{29*28*27*26*25}{5!} = ?\]